draw a cube and rectangular prism that would be filled exactly by 512 cubic yards of sand
Answer by edjones(7569)

factoring 512 gives 2^9
So (2^3)^3=512
2^3=8
Each side of the cube is 8 yards.
.
2^2*2^3*2^4=512
4*8*16 is one of the dimensions of a rectangular prism that is 512 yds.^3
.
Ed